GROWING OF METAL-MOLECULE CLUSTERS IN HELIUM DROPLETS

\underline{EVGENY LOGINOV}, LUIS GOMEZ and ANDREY VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089-0482.

The utility of a continuous beam of helium droplets for assembly of composite metal - molecular clusters is studied. Clusters of silver having up to about 10^{4} atoms and molecules such as methane, ethane, ethylene and acetylene with up to 10^{4} were obtained via sequential pickup of molecules by He droplets with average sizes in the range of 10^{5} to 10^{7} atoms. The IR spectra of the formed clusters showed a splitting of the C-H band, which we ascribed to molecules in the first layer in immediate contact with the surface of the metal clusters and to molecules in subsequent layers. Applications of this technique to a broader range of metal-molecule composites and surface deposition of formed clusters are discussed.